# DART: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

### v1

@@ 4.2 Latent Diffusion Noise Optimization-Based Control Using DART
Motion in-betweening aims to generate realistic motion frames that smoothly transition between a pair of history and goal keyframes. We consider a text-conditioned variant where an additional text prompt is inputted to specify the action semantics of the frames in between. We compare our method with DNO (Karunratanakul et al., 2024) and OmniControl (Xie et al., 2024). For a fair comparison, we train DART on the HML3D dataset same as the baselines. We evaluate using test sequences covering diverse actions, with the sequence lengths ranging from 2 to 4 seconds. The quantitative evaluations are shown in Tab. 3. We report the $L_{2}$ norm errors between the generated motion and the history motion and goal keyframe. We also evaluate the motion realism with the skate and jerk metrics. The skate metric (Ling et al., 2020; Zhang et al., 2018) calculates a scaled foot skating when in contact with the floor: $s=disp\cdot(2-2^{h/0.03})$, where $disp$ is the foot displacement in two consecutive frames, $h$ denotes the higher foot height in consecutive frames and 0.03m is the threshold value for contact. We do not calculate skate metric for sequences where the feet are not on a flat floor, such as crawling and climbing down stairs. Our method can generate the motions closest to the keyframe and show fewer skating and jerk artifacts. Our method effectively preserves the semantics specified by the text prompts, while the baseline DNO occasionally ignores the text prompts to reach the goal keyframe, as illustrated in the examples of pacing in circles and dancing in the supplementary video. This highlights the superior capability of our latent motion primitive-based DART in harmonizing spatial control and text semantic alignment.
Table 3: Quantitative evaluation of noise optimization-based In-between. The best results except dataset are in bold and ‘$\pm$’ indicates the 95% confidence interval.
@@
Dataset - - $2.83\pm 0.00$ $0.99\pm 0.00$
OmniControl $17.22\pm 0.80$ $5.88\pm 0.57$ $5.48\pm 0.90$ $1.26\pm 0.10$
DNO $2.15\pm 0.24$ $5.52\pm 1.86$ $5.12\pm 0.55$ $0.72\pm 0.07$
DART Opt.(ours) $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.54}\pm 0.03$ $\textbf{3.97}\pm 0.96$ $\textbf{0.71}\pm 0.03$
@@
We qualitatively show that our latent noise optimization control can be applied to human-scene interaction synthesis, where the goal is to control the human to interact naturally with the surrounding environment. Given the text prompts specifying the actions and durations, we use latent noise optimization to control the human to reach the goal joint location while adhering to the scene contact and collision constraints. The scenes are represented as signed distance fields for evaluating human-scene collision and contact constraints. We present generated interactions of climbing stairs and walking to sit on a chair in Fig. 3 and the supplementary video.
@@ 4.3 Reinforcent Learning-Based Control Using DART
Combining DART with reinforcement learning-based control, we train text-conditioned goal-reaching policies for three actions including ‘walk’, ‘run’, and ‘hop on the left leg’. We evaluate the controllers on paths consisting of sequences of waypoints. We compare our method to a baseline GAMMA (Zhang & Tang, 2022), which learns a motion model using the same motion source of AMASS (Mahmood et al., 2019) dataset as ours and trains RL controllers to reach goal locations. However, GAMMA does not support text conditioning and can only generate walking motion. The evaluation metrics include the reach time, the success rate of reaching the final goal waypoint, foot skating, and foot-floor distance. The evaluation results are shown in Tab. 4. Our controller consistently reaches all goals within a reasonable timeframe, while GAMMA occasionally fails to meet the final goal and may float off the floor beyond the contact threshold. Our text-conditioned goal-reaching policy achieves a generation speed of 240 frames per second. We believe our results on text-conditioned goal-reaching controllers demonstrate the potential of DART as a foundational human motion model, upon which control models for multiple tasks can be learned without collecting separate datasets for different tasks.
@@
Ours ‘walk’ $13.82\pm 0.10$ $\mathbf{1.0}\pm 0.0$ $5.07\pm 0.18$ $1.87\pm 0.02$
Ours ‘run’ $\mathbf{12.16}\pm 0.06$ $\mathbf{1.0}\pm 0.0$ $\textbf{4.70}\pm 0.13$ $2.02\pm 0.02$
Ours ‘hop on left leg’ $13.89\pm 0.12$ $\mathbf{1.0}\pm 0.0$ $41.18\pm 0.61$ $\mathbf{1.43}\pm 0.02$
@@ 5 Limitations and conclusions
Our method relies on motion sequences with frame-level aligned text annotations, as provided in BABEL, to achieve precise text-motion alignment and natural transition between actions. When trained on the coarse sentence-level motion labels from HML3D, the text-motion alignment degenerates for texts describing multiple actions, resulting in motions randomly switching between the described actions in a random order. This occurs because each short motion primitive inherently matches only a portion of the sequence’s semantics. Using a coarse sentence-level description as the text label for a primitive introduces semantic misalignment and ambiguity. Inspired by Stoffl et al. (2024), we aim to explore hierarchical latent spaces to effectively tackle both fine-grained and global sequence-level semantics in future work.
@@ Appendix A Motion Primitive Representation
Representation. We represent each frame of the motion primitive as a tuple of $(\mathbf{t},\mathbf{R},\mathbf{\theta},\mathbf{J},d\mathbf{t},d\mathbf{R},d\mathbf{J})$, where $\mathbf{t}\in\mathbb{R}^{3}$ denotes the global body translation, $\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D rotation representation (Zhou et al., 2020) of the global body orientation, $\mathbf{\theta}\in\mathbb{R}^{21\times 6}$ is the 6D representation of 21 joint rotations, $\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the 22 joints locations, $d\mathbf{t}\in\mathbb{R}^{3}$ denotes the temporal difference with previous frame’s translation, $d\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D representation of the relative rotation between current and previous frame’s body orientation, and $d\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the temporal diffence between current and previous frame’s joint locations.
Such overparameterized motion representation has multiple benefits. Firstly, the joint rotation components $\mathbf{\theta}$ are compatible with animation pipelines, saving the time-consuming optimization-based skeleton-to-body conversion required by the commonly used HML3D (Guo et al., 2022) representation in text-to-motion methods. Moreover, including the joint location components $\mathbf{J}$ facilitates solving physical constraints like reducing foot skating and joint trajectory control. Our motion representation also models the first-order kinematics with the temporal difference features to improve the motion naturalness.
@@ Appendix B Datasets
The BABEL dataset contains motion capture sequences with frame-aligned text labels that annotate the fine-grained semantics of actions. Fine-grained text labels in BABEL allow models to learn precise human action controls and natural transitions among actions. We use the motion data at a framerate of 30 frames per second the same as prior works (Barquero et al., 2024; Athanasiou et al., 2022). During training, motion primitives are randomly sampled from data sequences and the text label is randomly sampled from all the action segments that overlap with the primitive. To alleviate the action imbalance in the BABEL dataset, we use the provided action labels to perform importance sampling during training so that the motion data of each action has roughly equal sampling chances despite their varying frequency in the original dataset.
The HML3D dataset contains short motions with coarse sequence-level sentence descriptions. We only use a subset of HML3D in training since its subset HumanAct12 and left-right mirroring motion sequences only provide joint locations instead of SMPL (Loper et al., 2015; Pavlakos et al., 2019) body sequences that our motion representation requires. The joint rotations used in the original HML3D are calculated using naive inverse kinematics and can not be directly used to animate human motions. We therefore only train using the subset where the SMPL body motion sequences is available. The HML3D motions are sampled at a framerate of 20 frames per second. During training, we randomly sample primitives with uniform probability and the text label is randomly chosen from one of the multiple sentence captions of the overlapping sequence.
@@ Appendix C Motion Primitive VAE
Architecture. Our Motion Primitive VAE employs the transformer-based architecture. Both the encoder and decoder consist of 7 transformer encoder layers with skip connections (Chen et al., 2023). The transformer layers use the dropout rate of 0.1, feed-forward dimension of 1024, hidden dimension of 256, 4 attention heads, and the gelu activation function. The latent space dimension is 256. After finishing training, we follow (Rombach et al., 2022) to calculate the variance of the latent variables using a data batch. When training the latent denoiser model, the raw latent output from the encoder is scaled to have a unit standard deviation.
@@
$L_{VAE}=L_{rec}+w_{KL}\times L_{KL}+w_{aux}\times L_{aux}$ (5)
$L_{rec}=\mathcal{F}(\hat{\mathbf{X}},\mathbf{X})$ (6)
The reconstruction loss $L_{rec}$ aims to minimize the distance between the reconstructed future frames $\hat{\mathbf{X}}$ and the ground truth future frames $\mathbf{X}$, where $\mathcal{F}(\cdot,\cdot)$ denotes the distance function and we use the smoothed L1 loss(Girshick, 2015) in implementation.
$\mathcal{L}_{KL}=KL(q(z|\mathbf{H})||\mathcal{N}(0,I)),$ (7)
$L_{KL}$ penalizes the distribution difference between the predicted distribution and a standard Gaussian, where $KL(\cdot,\cdot)$ denotes the Kullback-Leibler divergence (KL), and $q(z|\mathbf{H})$ denotes the predicted distribution from the encoder $\mathcal{E}$. We use a small KL term of $1e^{-6}$ following Rombach et al. (2022) as we aim to keep the latent space expressive and only use the small KL loss to avoid arbitrarily high-variance latent spaces.
$L_{aux}=\mathcal{F}(\bar{d\mathbf{t}},\hat{d\mathbf{t}})+\mathcal{F}(\bar{d\mathbf{J}},\hat{d\mathbf{J}})+\mathcal{F}(\bar{d\mathbf{R}},\hat{d\mathbf{R}})$ (8)
The auxiliary loss $L_{aux}$ regularizes the predicted temporal difference features $\hat{d\cdot}$ of translations, global orientation, and joints to be close to the actual temporal differences $\bar{d\cdot}$ calculated from the predicted motions. We use $w_{KL}=1e^{-6}$ and $w_{aux}=100$ in our experiments. We train the motion primitive VAE with the AdamW (Kingma & Ba, 2017) optimizer and the learning rate is set to $1e^{-4}$ with linear annealing.
@@ D.1 Losses
$L_{simple}=\mathcal{F}(\mathcal{G}(z_{t},t,\mathbf{H},c),z_{0}),$ (10)
where $\mathcal{F}(\cdot,\cdot)$ is a distance function and we use the smooth L1 loss (Girshick, 2015) in our implementation. We train the denoiser to predict the clean latent variable with the simple objective $L_{simple}$, and apply the feature reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded motion primitive $\hat{\mathbf{X}}=\mathcal{D}(\mathcal{G}(z_{t},t,\mathbf{H},c))$ to ensure the decoded motion primitives are valid.

### v2

@@ 4.2 Latent Diffusion Noise Optimization-Based Control Using DART
Motion in-betweening aims to generate realistic motion frames that smoothly transition between a pair of history and goal keyframes. We consider a text-conditioned variant where an additional text prompt is inputted to specify the action semantics of the frames in between. We compare our method with DNO (Karunratanakul et al., 2024b) and OmniControl (Xie et al., 2024). For a fair comparison, we train DART on the HML3D dataset same as the baselines. We evaluate using test sequences covering diverse actions, with the sequence lengths ranging from 2 to 4 seconds. The quantitative evaluations are shown in Tab. 3. We report the $L_{2}$ norm errors between the generated motion and the history motion and goal keyframe. We also evaluate the motion realism with the skate and jerk metrics. The skate metric (Ling et al., 2020; Zhang et al., 2018) calculates a scaled foot skating when in contact with the floor: $s=disp\cdot(2-2^{h/0.03})$, where $disp$ is the foot displacement in two consecutive frames, $h$ denotes the higher foot height in consecutive frames and 0.03m is the threshold value for contact. We do not calculate skate metric for sequences where the feet are not on a flat floor, such as crawling and climbing down stairs. Our method can generate the motions closest to the keyframe and show fewer skating and jerk artifacts. Our method effectively preserves the semantics specified by the text prompts, while the baseline DNO occasionally ignores the text prompts to reach the goal keyframe, as illustrated in the examples of pacing in circles and dancing in the supplementary video. This highlights the superior capability of our latent motion primitive-based DART in harmonizing spatial control and text semantic alignment.
Table 3: Quantitative evaluation of noise optimization-based In-between. The best results excluding the dataset are in bold and ‘$\pm$’ indicates the 95% confidence interval.
@@
Dataset 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.27 $\pm$ 0.00 0.74 $\pm$ 0.00
OmniControl 21.22 $\pm$ 2.86 7.79 $\pm$ 1.91 4.97 $\pm$ 1.31 1.41 $\pm$ 0.08
DNO 1.20 $\pm$ 0.20 4.24 $\pm$ 1.34 5.38 $\pm$ 0.70 0.65 $\pm$ 0.06
Ours 0.00 $\pm$ 0.00 0.59 $\pm$ 0.01 2.98 $\pm$ 0.32 0.61 $\pm$ 0.01
@@
We qualitatively show that our latent noise optimization control can be applied to human-scene interaction synthesis, where the goal is to control the human to interact naturally with the surrounding environment. Given an input 3D scene and the text prompts specifying the actions and durations, we use latent noise optimization to control the human to reach the goal joint location while adhering to the scene contact and collision constraints. The input scenes are represented as signed distance fields for evaluating human-scene collision and contact constraints as detailed in Appendix F. We present generated interactions of climbing stairs and walking to sit on a chair in Fig. 3 and the supplementary video.
@@ 4.3 Reinforcent Learning-Based Control Using DART
By integrating DART with reinforcement learning-based control, we train text-conditioned goal-reaching policy models capable of three locomotion styles: ‘walk’, ‘run’, and ‘hop on the left leg’. We evaluate goal-reaching on paths consisting of sequences of waypoints. We compare our method to a baseline GAMMA (Zhang & Tang, 2022), which also trains goal-reaching policies with a learned motion action space. Unlike DART, GAMMA lacks text conditioning and is limited to generating walking motions. The evaluation metrics include the reach time, the success rate of reaching the final goal waypoint, foot skating, and foot-floor distance. The evaluation results are shown in Tab. 4. Our policy consistently reaches all goals within a reasonable timeframe, while GAMMA occasionally fails to meet the final goal and may float off the floor beyond the contact threshold. Moreover, our text-conditioned goal-reaching policy achieves a generation speed of 240 frames per second. These results demonstrate the potential of DART as a foundational human motion model, upon which versatile control models for various tasks can be learned through reinforcement learning.
@@
Ours ‘walk’ $17.08\pm 0.05$ $\mathbf{1.0}\pm 0.0$ $2.67\pm 0.12$ $\textbf{2.24}\pm 0.02$
Ours ‘run’ $\mathbf{10.55}\pm 0.06$ $\mathbf{1.0}\pm 0.0$ $3.23\pm 0.24$ $3.86\pm 0.05$
Ours ‘hop on left leg’ $20.50\pm 0.24$ $\mathbf{1.0}\pm 0.0$ $\textbf{2.22}\pm 0.12$ $4.11\pm 0.07$
@@ 5 Limitations and conclusions
DART relies on motion sequences with frame-level aligned text annotations, as in BABEL, to achieve precise text-motion alignment and natural transition between actions. When trained on the coarse sentence-level motion labels from HML3D, the text-motion alignment degenerates for texts describing multiple actions, resulting in motions randomly switching between the described actions in a random order. This occurs because each short motion primitive inherently matches only a portion of the sequence’s semantics. Using a coarse sentence-level description as the text label for a primitive causes semantic misalignment and ambiguity. We aim to explore hierarchical latent spaces to effectively tackle both fine-grained and global sequence-level semantics (Stoffl et al., 2024) in the future.
@@
Acknowledgements. We sincerely acknowledge the anonymous reviewers for their insightful feedback. We thank Korrawe Karunratanakul for helpful discussion and suggestions, and Siwei Zhang for meticulous proofreading. This work is supported by the SNSF project grant 200021 204840 and SDSC PhD fellowship.
@@ Appendix A Motion Primitive Representation
Representation. We represent each frame of the motion primitive as a tuple of $(\mathbf{t},\mathbf{R},\bm{\theta},\mathbf{J},\mathrm{d}\mathbf{t},\mathrm{d}\mathbf{R},\mathrm{d}\mathbf{J})$, where $\mathbf{t}\in\mathbb{R}^{3}$ denotes the global body translation, $\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D rotation representation (Zhou et al., 2019) of the global body orientation, $\bm{\theta}\in\mathbb{R}^{21\times 6}$ is the 6D representation of 21 joint rotations, $\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the 22 joints locations, $\mathrm{d}\mathbf{t}\in\mathbb{R}^{3}$ denotes the temporal difference with previous frame’s translation, $\mathrm{d}\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D representation of the relative rotation between current and previous frame’s body orientation, and $\mathrm{d}\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the temporal diffence between current and previous frame’s joint locations.
Our motion representation is overparameterized (Holden et al., 2017; Ling et al., 2020; Rempe et al., 2021; Guo et al., 2022), with multiple benefits. Firstly, the joint rotation components $\bm{\theta}$ are compatible with animation pipelines, saving the time-consuming optimization-based skeleton-to-body conversion required by the commonly used HML3D (Guo et al., 2022) representation in text-to-motion methods. Moreover, including the joint location components $\mathbf{J}$ facilitates solving physical constraints like reducing foot skating and joint trajectory control. Our motion representation also models the first-order kinematics with the temporal difference features to improve the motion naturalness.
@@ Appendix B Datasets
The BABEL dataset contains motion capture sequences with frame-aligned text labels that annotate the fine-grained semantics of actions. Fine-grained text labels in BABEL allow models to learn precise human action controls and natural transitions among actions. We use the motion data at a framerate of 30 frames per second the same as prior works (Barquero et al., 2024; Athanasiou et al., 2022). During training, motion primitives are randomly sampled from data sequences and the text label is randomly sampled from all the action segments that overlap with the primitive. To alleviate the action imbalance in the BABEL dataset, we use the provided action labels to perform importance sampling during training so that the motion data of each action category has roughly equal sampling chances despite their varying frequency in the original dataset. To maintain compatibility with prior works, we fix the human gender to male and set the body shape parameter to zero.
The HML3D dataset contains short motions with coarse sequence-level sentence descriptions. We only use a subset of HML3D in training since its subset HumanAct12 and left-right mirroring motion sequences only provide joint locations instead of SMPL (Loper et al., 2015; Pavlakos et al., 2019) body sequences that our motion representation requires. The joint rotations used in the original HML3D are calculated using naive inverse kinematics and can not be directly used to animate human motions. We therefore only train using the subset where the SMPL body motion sequences is available. The HML3D motions are sampled at a framerate of 20 frames per second. During training, we randomly sample primitives with uniform probability and the text label is randomly chosen from one of the multiple sentence captions of the overlapping sequence. To maintain compatibility with prior works, we fix the human gender to male and set the body shape parameter to zero.
@@ Appendix C Motion Primitive VAE
Architecture. Our Motion Primitive VAE employs the transformer-based architecture. Both the encoder and decoder consist of 7 transformer encoder layers with skip connections (Chen et al., 2023b). The transformer layers use the dropout rate of 0.1, feed-forward dimension of 1024, hidden dimension of 256, 4 attention heads, and the gelu activation function. The latent space dimension is 256. After finishing training, we follow (Rombach et al., 2022) to calculate the variance of the latent variables using a data batch. When training the latent denoiser model, the raw latent output from the encoder is scaled to have a unit standard deviation.
@@
$L_{VAE}=L_{rec}+w_{KL}\times L_{KL}+w_{aux}\times L_{aux}{\color[rgb]{0,0,0}+w_{SMPL}\times L_{SMPL}}.$ (4)
The reconstruction loss $L_{rec}$ aims to minimize the distance between the reconstructed future frames $\hat{\mathbf{X}}$ and the ground truth future frames $\mathbf{X}$ as follows:
$L_{rec}=\mathcal{F}(\hat{\mathbf{X}},\mathbf{X}),$ (5)
where $\mathcal{F}(\cdot,\cdot)$ denotes the distance function and we use the smoothed L1 loss(Girshick, 2015) in implementation.
The Kullback-Leibler divergence (KL) term $L_{KL}$ penalizes the distribution difference between the predicted distribution and a standard Gaussian:
$\mathcal{L}_{KL}=KL(q(\mathbf{z}|\mathbf{H})\parallel\mathcal{N}(0,\mathbf{I})),$ (6)
where $KL(\cdot,\cdot)$ denotes the Kullback-Leibler divergence (KL), and $q(\mathbf{z}|\mathbf{H})$ denotes the predicted distribution from the encoder $\mathcal{E}$. We use a small KL term of $1e^{-6}$ following Rombach et al. (2022) as we aim to keep the latent space expressive and only use the small KL loss to avoid arbitrarily high-variance latent spaces.
The auxiliary loss $L_{aux}$ regularizes the predicted temporal difference features $\hat{\mathrm{d}\cdot}$ of translations, global orientation, and joints to be close to the actual temporal differences $\bar{\mathrm{d}\cdot}$ calculated from the predicted motion features:
$L_{aux}=\mathcal{F}(\bar{\mathrm{d}\mathbf{t}},\hat{\mathrm{d}\mathbf{t}})+\mathcal{F}(\bar{\mathrm{d}\mathbf{J}},\hat{\mathrm{d}\mathbf{J}})+\mathcal{F}(\bar{\mathrm{d}\mathbf{R}},\hat{\mathrm{d}\mathbf{R}}).$ (7)
For instance, the difference of the first two frames of the predicted root translation ($\bar{\mathrm{d}\mathbf{t}}^{0}:=\hat{\mathbf{t}}^{1}-\hat{\mathbf{t}}^{0}$) should be consistent with the predicted first frame root translation difference feature $\hat{\mathrm{d}\mathbf{t}}^{0}$. We use $w_{aux}=100$ in our experiments.
The SMPL losses $L_{SMPL}$ consist of two components: the SMPL-based joint reconstruction loss $L_{joint\_rec}$, and the joint consistency loss $L_{consistency}$. The SMPL losses $L_{SMPL}$ are defined as follows:
${\color[rgb]{0,0,0}L_{SMPL}=L_{joint\_rec}+L_{consistency}.}$ (8)
The SMPL joint reconstruction loss $L_{joint\_rec}$ penalizes the discrepancy between the joints regressed from the predicted body parameters and those regressed from the ground truth body parameters:
${\color[rgb]{0,0,0}L_{joint\_rec}=\mathcal{F}(\mathcal{J}(\hat{\mathbf{t}},\hat{\mathbf{R}},\hat{\bm{\theta}}),\mathcal{J}(\mathbf{t},\mathbf{R},\bm{\theta})),}$ (9)
where $\mathcal{J}$ denotes the SMPL body joint regressor that predicts joint locations given the body parameters. The SMPL body shape parameter $\beta$ is assumed to be zero and ignored for simplicity here.
The joint consistency loss $L_{consistency}$ regularizes the predicted joint locations and predicted SMPL body parameters consistently represent the same body joints:
${\color[rgb]{0,0,0}L_{consistency}=\mathcal{F}(\hat{\mathbf{J}},\mathcal{J}(\hat{\mathbf{t}},\hat{\mathbf{R}},\hat{\bm{\theta}})).}$ (10)
We use $w_{SMPL}=100$ in our experiments. We train the motion primitive VAE with the AdamW (Kingma & Ba, 2015) optimizer and the learning rate is set to $1e^{-4}$ with linear annealing.
@@ D.1 Losses
$L_{simple}=\mathbb{E}_{(\mathbf{z}_{0},c)\sim q(\mathbf{z}_{0},c),\,t\sim[1,T],\,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\mathcal{F}(\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},c),\mathbf{z}_{0}),$ (12)
where $\mathcal{F}(\cdot,\cdot)$ is a distance function and we use the smooth L1 loss (Girshick, 2015) in our implementation. We train the denoiser to predict the clean latent variable with the simple objective $L_{simple}$, and apply the feature reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded motion primitive $\hat{\mathbf{X}}=\mathcal{D}(\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},c))$ to ensure the decoded motion primitives are valid.